OpenStudy (nicole365):

does anyone know how to do Logarithmic Equations and Inequalities? please help!!!!! i am soooo lost!!!!!

OpenStudy (anonymous):

name one, bet we can do it

OpenStudy (nicole365):

\[\log_{5} (-3x)<1\]

OpenStudy (anonymous):

ok we can do this

OpenStudy (nicole365):

thank u

OpenStudy (anonymous):

first thing to do is realize that log is an increasing function so not to worry

OpenStudy (anonymous):

do you know how to solve \[\log_5(-3x)=1\]? "no" is a fine answer, just asking

OpenStudy (nicole365):

no not really i was sick and didn't learn this at school

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] for example \[\log_5(25)=2\]is the same as \(5^2=25\)

OpenStudy (anonymous):

you need to know how to switch back and forth quickly \[2^4=16\iff \log_2(16)=4\]

OpenStudy (anonymous):

\[10^{-3}=0.001\iff \log_{10}(0.001)=-3\]

OpenStudy (anonymous):

\[\log_5(-3x)=1\iff -3x=5^1\]

OpenStudy (nicole365):

oh ok i get that thank u soo much it really helped

OpenStudy (anonymous):

now you have an extra step because you have an inequality, not an equality

OpenStudy (anonymous):

\[\log_5(-3x)<1\iff -3x<5^1\] o rmore simply \[-3x<5\] do you know how to solve that ?

OpenStudy (nicole365):

yes thank u!!! u made it sound and look soo easy i think i got it now!!!!!!

OpenStudy (anonymous):

make sure to switch the inequality when you divide by \(-3\) just to make sure, want to try another?

OpenStudy (nicole365):

k i wont!! nah i think i can do the others now i only have 3 more problems

OpenStudy (anonymous):

ok if you have questions let me know